# The $J/\psi\to\phi\eta\pi^{0}$ reaction: $a_{0}(980)$ production, triangle singularity and non $\phi$ background

### v1

@@ II.4 Tree level with other $K^{*}$ intermediate states
In Fig. 7 we show the $\phi\eta$ mass distribution. As discussed above, we have fitted the two peaks at around $1650\textrm{ MeV}$ and $1850\textrm{ MeV}$ with Breit Wigner distributions, so we must pay attention to the other structure appearing around $2400\textrm{ MeV}$, as a broad peak, that in principle should come from the mechanisms that we have used to describe the data. We see indeed, that a large fraction of the strength of this peak comes from the tree level contributions, which also produce a small peak around 2800 MeV, as seen in the experiment. We also see that the tree level contribution produces a broad bump around 2400 MeV where the experimental peak appears. However, this bump is not sufficient to reproduce the strength of the experimental peak. It is not clear what the reason for this discrepancy can be, or whether it is the signal of some resonance. However, the non $\phi$ contribution of the experimental analysis of Ref. [28], reproduces the peak, and we see that the missing strength in this region is reflected in the $\phi\pi^{0}$ mass distribution of Fig. 4 in missing strength around $1700\textrm{ MeV}$ between the two peaks coming from the tree level contributions with intermediate $K^{*}(890)$ and $K^{*}(1410)$ resonances. Hence, most likely it reflects the shape of $K^{*}(890)$ and $K^{*}(1410)$ resonances obtained assuming propagators with a constant width. But whatever the reason, we tend to think that this is not a signal of a new state. It could also be affected by effects on the $h_{1}(1415)$ resonance (before called $h_{1}(1380)$) from reinteraction of $K^{*}\bar{K}$ in Fig. 3 (see Ref. [53]). These effects at the peak of the $\phi\pi^{0}$ mass distribution around $1400\textrm{ MeV}$ are effectively incorporated in our $A_{tT,890}$ constant and are the same in the triangle diagram of Fig. 2 at the peak of the triangle singularity. Hence the ratio of the strength of these two peaks is well calculated, but there would be corrections in the tree level contribution at higher invariant mass which would modify the shape at Fig. 4 in the dip region.
@@
At this point it is interesting to discuss why the peak of the TS and the non $\phi$ distributions associated to the $K^{*}(890)$ tree level contribution appear at the same energy in the $M_{\phi\pi^{0}}$ distribution. Indeed, as we discussed, a TS in the diagrams of Fig. 2 appears when all the three intermediate particles are on shell and are collinear. This situation will be repeated in the diagrams of Fig. 3, since the $K^{+}$ and $K^{-}$ are external particles and hence, on shell. And in this case the $K^{*}$ propagator will be placed on shell, and have its maximum, even being infinite if the width is neglected. The argument can also be repeated for the contribution of the $K^{*}(1410)$. In this case, related to the peak observed from the tree level $K^{+}K^{-}$ production term, one also finds a TS from the diagrams of Fig. 2, with the substitution $K^{*}\to K^{*}(1410)$, peaking at the same place as the tree level but with a much smaller strength.
@@
It has become clear that the non $\phi$ contribution in this reaction blurs the signals of the $TS$. This is due to the experimental mode used in Ref. [28] to identify the $\phi$, by looking at its $K^{+}K^{-}$ decay. If one identified the $\phi$ with other method or other decay mode, say three pions for instance, this undesired background would disappear, allowing one to observe the TSs that are unavoidable, since their strength is tied to the non $\phi$ peak observed in Ref. [28] that we have fitted, and thus our predictions are reliable. A different way to observe the TS related to this process is by looking at the $J/\psi\to\phi\pi^{+}a_{0}(980)$ reaction, dicussed in Ref. [54].
@@
IV Conclusions
@@
We have carried out an analysis of the $J/\psi\to\phi\eta\pi^{0}$ reaction recently measured with high precision by the BESIII collaboration [28], where the $a_{0}(980)$ resonance, which appears in an isospin forbidden mode, is clearly seen. Explaining this production mode is one of the purposes of the present work. Another motivation was the prediction in Ref. [27] that there should be a triangle singularity from a reaction mechanism, showing a peak in the $\phi\pi^{0}$ mass distribution around $1385\textrm{ MeV}$ [27]. The experimental data show a clear peak around this energy, which however was identified in the experimental analysis as a “non $\phi$” contribution. Trying to decipher the origin of this “non $\phi$” contribution is also an issue that we tackle in the present work.
@@
We also found the origin of the peaks in the $\phi\pi^{0}$ mass distribution as coming from the way the $\phi$ was identified in the experiment measuring $K^{+}K^{-}$ in a narrow window around the $\phi$ mass. However, we could show that there was a tree level mechanism for $J/\psi\to\eta K^{+}K^{-}\pi^{0}$, where the pair $K^{+},K^{-}$ was produced in isospin $I=1$, hence not coming from $\phi$ decay, and provided a contribution to the mass distribution. The mechanism for this tree level contribution is $J/\psi\to\eta K^{-}K^{*+};\,K^{*+}\to\pi^{0}K^{+}$ plus $J/\psi\to\eta K^{+}K^{*-};\,K^{*-}\to\pi^{0}K^{-}$. If the intermediate $K^{*}$ is $K^{*}(890)$ a peak appears around $1390\textrm{ MeV}$, and if the intermediate $K^{*}$ is $K^{*}(1410)$ a peak appears around $2050\textrm{ MeV}$, and both peaks are seen in the experimental data.
Related to these latter mechanism there is a triangle mechanism, suggested in Ref. [27], where the $K^{+}$ and $K^{-}$ merge to produce $\pi^{0}\eta$ and hence the $a_{0}(980)$. The strength of this mechanism is tied to that of the tree level one, since only the $K\bar{K}\to\pi^{0}\eta$ transition amplitudes, well known theoretically, are needed to complete the loop diagram. Then we can make a good prediction of the strength of this mode, also violating isospin, which turns out to be about a factor $40$ smaller than the peaks originated from the tree level mechanisms that do not involve isospin violation.
@@
In summary, we provided an explanation for the $a_{0}(980)$ production mode, with the abnormally narrow width observed in the experiment. We also could identify the origin of the peaks in the $\phi\pi^{0}$ mass distribution observed in the experiment, branded as “non $\phi$” contribution in the experimental analysis, and finally we could see that there are indeed two peaks in the $\phi\pi^{0}$ mass distribution showing a triangle singularity, but their strength is significantly smaller than the observed peaks. We also mentioned that to see these peaks the technique used so far to identify the $\phi$, by looking at its $K^{+}K^{-}$ decay, would have to be changed to avoid contributions from $K^{+}K^{-}$ that do not come from $\phi$ decay and have a large strength. The results obtained here should be of use for planning future experiments aiming at identifying the triangle singularities predicted here.
@@ Acknowledgments
Appendix A Four body phase space calculation and invariant mass distributions
The decay width for $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ is given by
@@ Appendix A Four body phase space calculation and invariant mass distributions
$\Theta(1-A^{2}_{\theta})\;\Theta(M_{J/\psi}-E_{\eta}-E_{K^{+}}-E_{K^{-}}),$ (42)
@@
The rest of the work is done using standard Monte Carlo calculations generating events over an enlarged phase space that contains the actual phase space and impose the $K^{+}K^{-}$ cut of Eq. (21). For every good MC event, different invariant masses are evaluated and stored in boxes, which at the end provide the differential mass distributions. Apart from the trivial limits of the different variables, it is worth mentioning the maximum values of $P_{K}$ and $q_{K}$. The former corresponds to the case when the $K^{+}K^{-}$ and $\pi^{0}\eta$ pairs have their minimum invariant masses. Thus,

### v2

@@ II.4 Tree level with other $K^{*}$ intermediate states
We show the $\phi\eta$ mass distribution in Fig. 8 and only the TS contribution in Fig. 9. As discussed above, we have fitted the two peaks at around $1650$ and $1850\textrm{ MeV}$ with Breit-Wigner distributions, so we must pay attention to the other structure appearing around $2400\textrm{ MeV}$, as a broad peak, that, in principle, should come from the mechanisms that we have used to describe the data. We see indeed, that a large fraction of the strength of this peak comes from the tree level contributions, which also produce a small peak around 2800 MeV, as seen in the experiment. We also see that the tree level contribution produces a broad bump around 2400 MeV where the experimental peak appears. However, this bump is not sufficient to reproduce the strength of the experimental peak. It is not clear what the reason for this discrepancy can be, or whether it is the signal of some resonance. However, the non-$\phi$ contribution of the experimental analysis of Ref. [28], reproduces the peak, and we see that the missing strength in this region is reflected in the $\phi\pi^{0}$ mass distribution of Fig. 4 in missing strength around $1700\textrm{ MeV}$ between the two peaks coming from the tree level contributions with intermediate $K^{*}(890)$ and $K^{*}(1410)$ resonances. Hence, most likely it reflects the shape of $K^{*}(890)$ and $K^{*}(1410)$ resonances obtained assuming propagators with a constant width. Whatever the reason, we tend to think that this is not a signal of a new state. It could also be affected by effects on the $h_{1}(1415)$ resonance [before called $h_{1}(1380)$] from reinteraction of $K^{*}\bar{K}$ in Fig. 3 (see Ref. [53]). These effects at the peak of the $\phi\pi^{0}$ mass distribution around $1400\textrm{ MeV}$ are effectively incorporated in our $A_{tT,890}$ constant and are the same in the triangle diagram of Fig. 2 at the peak of the triangle singularity. Hence the ratio of the strength of these two peaks is well calculated, but there would be corrections in the tree level contribution at higher invariant mass that would modify the shape at Fig. 4 in the dip region.
@@
At this point it is interesting to discuss why the peak of the TS and the non-$\phi$ distributions associated with the $K^{*}(890)$ tree level contribution appear at the same energy in the $M_{\phi\pi^{0}}$ distribution. Indeed, as we discussed, a TS in the diagrams of Fig. 2 appears when all the three intermediate particles are on shell and are collinear. This situation will be repeated in the diagrams of Fig. 3, since the $K^{+}$ and $K^{-}$ are external particles and hence, on shell. And in this case the $K^{*}$ propagator will be placed on shell, and have its maximum, even being infinite if the width is neglected. The argument can also be repeated for the contribution of the $K^{*}(1410)$. In this case, related to the peak observed from the tree level $K^{+}K^{-}$ production term, one also finds a TS from the diagrams of Fig. 2, with the substitution $K^{*}\to K^{*}(1410)$, peaking at the same place as the tree level but with a much smaller strength.
@@
It has become clear that the non-$\phi$ contribution in this reaction blurs the signals of the TS. This is due to the experimental mode used in Ref. [28] to identify the $\phi$, by looking at its $K^{+}K^{-}$ decay. If one identified the $\phi$ with another method or other decay mode, say three pions for instance, this undesired background would disappear, allowing one to observe the TSs that are unavoidable, since their strength is tied to the non-$\phi$ peak observed in Ref. [28] that we have fitted, and thus our predictions are reliable. A different way to observe the TS related to this process is by looking at the $J/\psi\to\phi\pi^{+}a_{0}(980)$ reaction, discussed in Ref. [54].
@@
IV FURTHER INTERACTION STEPS
We have $J/\psi$ decay to four particles $K^{+}K^{-}\eta\pi^{0}$ and we have considered interactions of pairs. In Fig. 1 we have $K^{+}K^{-}$ coming from the $\phi$ decay, and $K\bar{K}\to\pi^{0}\eta$. In Fig. 3 we have $K^{*}\to\pi K$. In the triangle diagrams of Fig. 2 we have a three-body interaction, $K^{*}\to\pi K$ and the $K$ coming from this decay interacts with a third particle. We found that this mechanism, even producing a triangle singularity, gives a very small signal.
In the case of three particles in the final state, there is the traditional approach of Khuri-Treiman [55, 56] which accounts for multiple scattering of particles, accounting for the interaction of a pair followed by reinteraction of one of these particles with the third particle. Our triangle diagrams of Fig. 2 would be one example of interaction beyond the interaction of just pairs. The Khuri-Treiman approach is of much value when studying the decay of light particles, where there is little phase space and the kinematical conditions in rescattering are not very different from those of the first scattering. Such is the case in $\eta\to 3\pi$ [57], the $\omega/\phi\to 3\pi$ [58] or the $\phi\to 3\pi$ in Ref. [59]. In decays of heavy hadrons, there is more phase space and the scattering of a pair normally selects some resonance. Fixing the two-body kinematics to this resonance, the range of invariant mass of one of the particles of the pair with the third particle has limitations that make it unlikely to match with another resonance for the new pair, the only way to make competitive the new mechanism with an extra loop. This has been discussed in detail in Sec. IV A of Ref. [60] in the $\Lambda_{b}\to D^{+}D^{-}\Lambda$, $D^{0}D_{s}^{-}p$, $D_{s}^{+}D_{s}^{-}\Lambda$ reactions. Yet, we find it instructive to discuss the consequence of going one step further in the interaction in the mechanisms that we have.
Let us start from the mechanism of Fig. 1, where the $K\bar{K}\to\pi^{0}\eta$ interaction proceeds through the excitation of the $a_{0}(980)$ resonance, as shown in Fig. 6. We can depict the rescattering mechanisms as shown in Fig. 10.
Figure 10: Rescattering mechanisms from Fig. 1: (a) $\phi\pi^{0}$ interaction generating the $b_{1}(1235)$ resonance; (b) $\phi\eta$ interaction generating the $h_{1}(1170)$, $h_{1}(1415)$ resonances.
These are triangle diagrams, which could develop or not a triangle singularity. From the experience of the results from the triangle diagrams of Fig. 2, we might expect a small contribution with respect to the large ones evaluated before, even more if the diagrams of Fig. 10 do not develop a TS.
To estimate the relevance of the diagrams of Fig. 10, we use two sources of information:
  • 1)
    Which resonances couple to $\phi\pi^{0}$ and $\phi\eta$? We find this information in the work of Ref. [53].
  • 2)
    How close or far away are we from having a TS in these diagrams? We find this information in Eq. (18) of Ref. [39], which is the test that all particles in the loop are on shell and collinear, satisfying the Coleman-Norton theorem [38] that the mechanism can proceed at the classical level.
In Ref. [53] we find that the $\phi\pi^{0}$ channel couples strongly to the $b_{1}(1235)$ resonance. Direct application of Eq. (18) of Ref. [39] to the diagram of Fig. 10(a) tells us that a TS develops if the mass of the $J/\psi$ was $2027\textrm{ MeV}$, very far away from the actual mass of the $J/\psi$, which makes that mechanism negligible.
Concerning the diagram of Fig. 10(b), the $\phi\eta$ channel couples to the $h_{1}$ resonances, $h_{1}(1170)$ and $h_{1}(1380)$ [now $h_{1}(1415)$]. However, as found in Ref. [53], the coupling of $\phi\eta$ to the $h_{1}(1170)$ is extremely small, and hence only the $h_{1}(1415)$ that has an appreciable coupling to $\phi\eta$ could be relevant. Once again, application of Eq. (18) of Ref. [39] tells us that a TS develops if the mass of the $J/\psi$ was $2049\textrm{ MeV}$, far away from the actual mass, making again the contribution of this mechanism negligible.
Figure 11: Rescattering mechanisms of Fig. 3 through $\eta K$ interaction.
Next we turn our attention to the important diagrams of Fig. 3 and proceed with possible rescatterings. The $K\bar{K}$ rescattering leads to the triangle mechanisms of Fig. 2, which develop a TS, in spite of which, the contribution is very small. Then we consider the $K\eta$ interaction. This leads us to the mechanism depicted in Fig. 11. This would add a term with the former amplitude multiplied by $G_{K\eta}\;t_{K\eta,\,K\eta}$. The $K\eta$ couples to the $K_{0}^{*}(700)$ resonance, but its coupling is smaller than to $\pi K$, and appears squared in $t_{K\eta,\,K\eta}$, rendering a factor of about 3 reduction with respect to $t_{\pi K,\,\pi K}$ [61], but more important, $m_{K}+m_{\eta}=1041\textrm{ MeV}$, far away from the nominal mass of the $K_{0}^{*}(700)$ even considering its large width. Hence, this contribution is also very small.
Next we make the $K^{+}$ or the $\pi^{0}$ of Fig. 3(a) reinteract with the $\eta$ or $K^{-}$ and we get three new mechanisms depicted in Fig. 12.
Figure 12: Rescattering mechanisms from Fig. 3 leading to triangle diagrams: (a) with $\pi\eta$ rescattering; (b) with $\pi\bar{K}$ rescattering; (c) with $\eta K$ rescattering.
The diagram of Fig. 12(a) would produce the $\eta\pi$ with the $a_{0}(980)$ resonance. Once again, applying Eq. (18) of Ref. [39], we find that one is extremely far away from satisfying the conditions of a TS and, again, the mechanism is totally negligible. In Fig. 12(c), the $\eta K$ couples to the $K^{*}_{0}(700)$ but it cannot develop a TS because the $\eta K$ mass exceeds the one of the $K^{*}_{0}(700)$ by far, even considering the width; in addition, as already discussed above, the amplitude $\eta K\to\eta K$ through the resonance is very small for the actual energies of $\eta$ and $K$ in the phase space.
[…]
V Conclusions
@@
We have carried out an analysis of the $J/\psi\to\phi\eta\pi^{0}$ reaction recently measured with high precision by the BESIII Collaboration [28], where the $a_{0}(980)$ resonance, which appears in an isospin forbidden mode, is clearly seen. Explaining this production mode is one of the purposes of the present work. Another motivation was the prediction in Ref. [27] that there should be a triangle singularity from a reaction mechanism, showing a peak in the $\phi\pi^{0}$ mass distribution around $1385\textrm{ MeV}$ [27]. The experimental data show a clear peak around this energy, which, however, was identified in the experimental analysis as a “non-$\phi$” contribution. Trying to decipher the origin of this “non-$\phi$” contribution is also an issue that we tackle in the present work.
@@
We also found the origin of the peaks in the $\phi\pi^{0}$ mass distribution as coming from the way the $\phi$ was identified in the experiment measuring $K^{+}K^{-}$ in a narrow window around the $\phi$ mass. However, we could show that there was a tree level mechanism for $J/\psi\to\eta K^{+}K^{-}\pi^{0}$, where the pair $K^{+},K^{-}$ was produced in isospin $I=1$, hence not coming from $\phi$ decay, and provided a contribution to the mass distribution. The mechanism for this tree level contribution is $J/\psi\to\eta K^{-}K^{*+};\,K^{*+}\to\pi^{0}K^{+}$ plus $J/\psi\to\eta K^{+}K^{*-};\,K^{*-}\to\pi^{0}K^{-}$. If the intermediate $K^{*}$ is $K^{*}(890)$, a peak appears around $1390\textrm{ MeV}$, and if the intermediate $K^{*}$ is $K^{*}(1410)$, a peak appears around $2050\textrm{ MeV}$, and both peaks are seen in the experimental data.
Related to these latter mechanisms there is a triangle mechanism, suggested in Ref. [27], where the $K^{+}$ and $K^{-}$ merge to produce $\pi^{0}\eta$ and hence the $a_{0}(980)$. The strength of this mechanism is tied to that of the tree level one, since only the $K\bar{K}\to\pi^{0}\eta$ transition amplitudes, well known theoretically, are needed to complete the loop diagram. Then we can make a good prediction of the strength of this mode, also violating isospin, which turns out to be about a factor $40$ smaller than the peaks that originated from the tree level mechanisms that do not involve isospin violation.
@@
In addition, we made a thorough discussion of possible effects of extra interactions beyond two-body interactions and found the contributions negligible.
In summary, we provided an explanation for the $a_{0}(980)$ production mode, with the abnormally narrow width observed in the experiment. We also could identify the origin of the peaks in the $\phi\pi^{0}$ mass distribution observed in the experiment, branded as “non-$\phi$” contribution in the experimental analysis; finally we could see that there are indeed two peaks in the $\phi\pi^{0}$ mass distribution showing a triangle singularity, but their strength is significantly smaller than that of the observed peaks. We also mentioned that to see these peaks the technique used so far to identify the $\phi$, by looking at its $K^{+}K^{-}$ decay, would have to be changed to avoid contributions from $K^{+}K^{-}$ that do not come from $\phi$ decay and have a large strength. The results obtained here should be of use for planning future experiments aiming at identifying the triangle singularities predicted here.
@@ Acknowledgments
Appendix A Four-body phase space calculation and invariant mass distributions
We follow the step of Ref. [62]. The decay width for $J/\psi\to\pi^{0}\eta K^{+}K^{-}$ is given by
@@ Appendix A Four-body phase space calculation and invariant mass distributions
$\Theta(1-A^{2}_{\theta})\;\Theta(M_{J/\psi}-E_{\eta}-E_{K^{+}}-E_{K^{-}}).$ (42)
@@
The rest of the work is done using standard Monte Carlo (MC) calculations generating events over an enlarged phase space that contains the actual phase space and impose the $K^{+}K^{-}$ cut of Eq. (21). For every good MC event, different invariant masses are evaluated and stored in boxes, which at the end provide the differential mass distributions. Apart from the trivial limits of the different variables, it is worth mentioning the maximum values of $P_{K}$ and $q_{K}$. The former corresponds to the case when the $K^{+}K^{-}$ and $\pi^{0}\eta$ pairs have their minimum invariant masses. Thus,